OpenStudy (anonymous):

find absolute minimum at (x,y) f(x) = 2 (square root x) (x − 1); x ≥ 0

OpenStudy (carson889):

2*x^(1/2) * (x-1) = 2*x^(3/2) - 2*x^(1/2) Take derivative: (6/2)(x^(1/2)) - x^(-1/2) Set equal to zero to find critical point: (6/2)(x^(1/2)) - x^(-1/2) = 0 (6/2)(x^(1/2)) = x^(-1/2) \[\frac{ 6\sqrt{x} }{ 2 }=\frac{ 1 }{ \sqrt{x} }\] Multiply both sides by sqrt(x): \[\frac{ 6x }{ 2 } = 1\] \[6x = 2\] \[x = \frac{ 1 }{ 3 }\] Plug x into original equation to get: \[\frac{ -4\sqrt{1/3} }{ 3 }\]

OpenStudy (anonymous):

@carson889 i just plugged that into my assignment and it stated it was incorrect. did you write down the problem correctly?

OpenStudy (anonymous):

\[2\sqrt{x}(x-1) ; x \ge0\]

OpenStudy (anonymous):

nvm i got it. thanks for your help

OpenStudy (carson889):

Your welcome. And glad it worked out.